# Reconstruction of Differentially Private Text Sanitization via Large Language Models

### v2

@@ 5.2 Datasets
Table 8: Results of the black-box instruction-based attacks against the word-level DP on WikiMIA.
@@
Table 9: Results of the black-box instruction-based attacks against the sentence-level DP on WikiMIA.
@@
Table 10: Results of the white-box fine-tuning-based attacks against the word-level DP on Pile-CC.
@@
Table 11: Results of the white-box fine-tuning-based attacks against the word-level DP on Pile-Wiki.
@@
Table 12: Results of the white-box fine-tuning-based attacks against the word-level DP on Pile-Enron.
@@
Table 13: Example of the white-box fine-tuning attacks against the word-level DP on Pile-CC and Pile-Enron.
Pile-CC Pile-Enron
@@
5.3 Differentially Private Text Sanitization Approaches
In the experiments, we choose MadLib [25] (Algorithm 2) as the word-level DP in our experiments, since this is the commonly used DP word-level implementation [41, 20]. In MadLib, the distance function is defined in Euclidean space and Laplace noise is added to the word embedding vector. We choose DP-Prompt [57] (Algorithm 3) as the sentence-level DP in our experiments, since Utpala et al. [57] demonstrate the DP-Prompt achieves better utility compared to other approaches when using ChatGPT. In DP-Prompt, we consider using ChatGPT-3.5 as base model to generate the paraphrased text.
5.4 Metrics
@@
  • Recall and Precision. In our study, we focus on the privacy leakage of sensitive content in the text rather than the shared vocabulary (e.g., time, addresses, and names), which are more prone to causing privacy concerns; whereas common words like “the” and “this” do not lead to privacy issues. That is, we aim to recover the privacy-sensitive content from the original text. Hence, we extract and mark PII sequences (see Section 2 for details) from the text and consider these sequences as the privacy within the given text. Formally, we denote the set of PII sequences in the original text $\boldsymbol{x}$ as $C$, in the DP-sanitized text $\tilde{\boldsymbol{x}}$ as $\tilde{C}$, and in the reconstructed text $\hat{\boldsymbol{x}}$ as $\hat{C}$. We define recall and precision for the reconstruction attacks below.
@@ 5.5 Implementation
To tag PII sequences from text, we leverage Flair [10] as the PII extraction model released in HuggingFace [2]. Specifically, Flair is a NER framework based on Transformer networks to classify tokens as PII sequences. It is noteworthy that in our experiments we did not store any PII as auxiliary knowledge for our attacks. For the models giving weight access (e.g. LLaMA, Gemma and OPT), we downloaded checkpoints from HuggingFace. For the models not giving weight access (e.g. ChatGPT, Claude), we query models via black-box APIs available online (e.g., OpenAI API [4] and Anthropic AI API [1]). For DP text sanitization approaches, we implement both the word-level DP and the sentence-level DP based on DP-Prompt repository [56]. Note that we implement the word-level DP using 50-dimensional GloVe [49] vectors (following existing works [41, 57]) as word embeddings. We implement our reconstruction attacks based on PyTorch [47] and Transformers [61]. To fine-tune the pre-trained LLMs, we leverage LoRA [29] based on PEFT [39] library, a parameter-efficient fine-tuning method, to reduce the trainable parameters due to the limited computational resource. To obtain the sensitive information similarity by LLMs, we employ ChatGPT-3.5 as the base model. Table 5 presents the prompt templates used for evaluating the sensitive information similarity. Finally, we run all the experiments on a GPU cluster, which is equipped with NVIDIA RTX3090, RTX4090 and A800.
5.6 Results of Black-box Instruction-based Attacks
@@
Setup. To evaluate the effectiveness of the black-box instruction-based attack, we conduct experiments on WikiMIA using both the word-level DP and the sentence-level DP under various privacy budgets ($\epsilon=\{4,8,12\}$ for the word-level DP [24] or $T=\{2.0,1.5,1.0\}$ for the sentence level DP [57], where the relationship between $T$ and the privacy budget can be found in Section 2) with instruction-tuned LLMs, including LLaMA, Gemma, ChatGPT, and Claude. We adjust the prompt per Figure 2 for different models and query models with the sanitized text.
@@
Results. Table 6 and Table 7 provides an example of the black-box instruction-based attack on WikiMIA under $\epsilon=8$ using ChatGPT-4, Claude-3-Opus, LLaMA-3-70B and Gemma-2-27B for the word-level DP and the sentence-level DP, repectively. The original text $\boldsymbol{x}$, sanitized text $\tilde{\boldsymbol{x}}$ and reconstructed text $\hat{\boldsymbol{x}}$ are listed, where the highlighted text is PII sequences extracted by the NER model, and the text in the red box contains the PII sequences that remain unchanged. For word-level DP, we observe that the reconstructed text can correct the errors in the sanitized text (e.g., reconstructed text using ChatGPT-4 correct “31st” to “51st”). Compared to word-level DP, sentence-level DP retains most of the PII sequences from the original text, thereby fewer PII sequences are successfully reconstructed.
@@
The numeric results on WikiMIA for the word-level DP are detailed in Table 8 and Table 9, respectively. For the word-level DP, in a practical privacy budgets (e.g., $\epsilon=8$ and $\epsilon=12$), the black-box instruction-based attacks successfully reconstruct sensitive information at the word level with high probability. Nevertheless, the black-box instruction-based attacks have no significant effect under a smaller privacy budget due to substantial discrepancies between the original and sanitized texts. For the sentence-level DP, we observe that Succ, Recall, and Precision are lower than the word-level DP, but Score is higher. The reason behind it is in two folds. First, the sentence-level DP fails to remove the sensitive information (see examples as the text enclosed in the red boxes in Table 9). Second, the metrics of recall, precision, and succ measures the difference between the DP-sanitized text and the reconstructed text. Hence, these three metrics do not show a promising results.
@@
5.7 Results of White-box Fine-tuning-based Attacks
@@
Results. Table 13 provides an example of the white-box fine-tuning-based attacks on Pile-CC under $\epsilon=8$ with Pythia-6.9B, OPT-6.7B and GPT-J-6B for the word-level DP. We observe that, the open-source model reconstruct few PII sequences compared to close-source models (e.g., ChatGPT, Claude).
@@
Table 10, Table 11, and Table 12 report the reconstruction results and Precision on Pile-CC, Pile-Wiki and Pile-Enron, respectively, where the metrics include Succ, Recall and Precision. Similar to the black-box instruction-based attacks, we observe that the model with more parameters can achieve higher performance metrics. For Pile-CC, the metrics including Succ, Recall, and Precision are lower than Pile-Wiki and Pile-Entron, this may be because Pile-CC is not composed of a specific type of data. Therefore, the model cannot learn from a specific data distribution like fact (Pile-Wiki) or email (Pile-Enron) distribution.
6 Related Work
Comparision with existing studies. Existing works discovered the "current" privacy concerns, which could be addresses by DP, the de facto privacy standard. However, our work, focusing on "future" privacy concerns as motivated in Section 1, demonstrates that DP might have the potential vulnerability in the era of LLMs, which can recover the privacy from the DP-sanitized text.
7 Mitigation & Discussion
PII classes. We also conducted experiments to show the performance of our attacks for different PII classes, which have different degrees of sensitive information leakage. Direct identifiers (e.g., names) leak more sensitive information than quasi-identifiers (e.g., addresses). For experimental settings, we select six classes (e.g., person, address, organization), and apply the black-box instruction-based attack to ChatGPT-4 with the word-level DP under privacy budget $\epsilon=8.0$. Figure 3 in Appendix F shows the precision and recall for different PII classes on WikiMIA. The results demonstrate that our reconstruction attacks exhibit varying levels of effectiveness in recovering different types of sensitive information.
Limitation. In this study, we could not conduct useful ablation studies to determine whether the training data played a crucial role in LLMs’ ability to recover privacy. On one hand, performing such ablation studies on black-box attacks is infeasible due to the unpublished training data of the LLMs, despite their promising recovery rates. On the other hand, for LLMs with specified training data, we tested white-box attacks using DP-sanitized text, where the corresponding original text was not included in the LLMs’ training data. Specifically, we prompted the same fine-tuned LLMs as those in Table 11 using DP-sanitized text from WikiMIA (which was guaranteed not to have been seen by the LLMs). The results, shown in Table 16 in Appendix F, indicate similar recovery rates to those in Table 11. However, these similar results further confirm that white-box LLMs trained on a limited number of text data do not exhibit sufficient capability to perform our attacks.
8 Conclusion
Ethical Considerations
Specifically, this paper discovered the capability of the modern large language models (LLMs) on recovering private content from differentially private prompts where the private information was sanitized. Such findings were based on experiments over the publicly available data and the open APIs to publicly available LLMs (see details in Section 5). We faithfully followed the Terms of Service of LLMs when interacting with the LLMs. Since it was believed that differentially private outputs were not invertible, our work might have the following negative outcomes for public interests.
  • Disclosure. We discovered a vulnerability in the existing privacy-preserving technique of differential privacy. According to the paper, adversaries could use the latest implementations of LLMs to recover historical data that was intended to be protected by differential privacy. To mitigate this vulnerability, we discussed potential solutions in Section 7.
We did not engage in any malicious activities, such as exposing sensitive information, disrupting legitimate services, or causing financial or reputational harm to the LLM vendors providing these services.
Compliance with the Open Science Policy
This paper adheres to the principles of open science by ensuring that all research data and complete source code used in the study will be publicly available upon the acceptance of the paper to foster further research.
Appendix
Appendix A Notations
Table 14 gives the summary of notations used in this paper.
Table 14: Summary of Notations.
Appendix B Language Models
Appendix C Word- and Sentence-level DP Approaches
Word-level DP approaches. The goal of the word-level DP is to perturb each word in a sentence under the metric DP. Formally, consider a sequence $\boldsymbol{x}=\{x_{1},x_{2},\dots,x_{n}\}$ with $n$ tokens, each token $x_{i}$ is convert to a $m$-dimensional word embedding $\phi_{i}=\phi(x_{i})$ by a pre-trained word embedding model $\phi:\mathcal{V}\rightarrow\mathbb{R}^{m}$, where $\mathcal{V}$ represents the vocabulary. To obtain the noisy embedding $\tilde{\phi}_{i}$, a noise $\boldsymbol{z}$ sampled from a multivariate probability distribution $p_{\epsilon}(\boldsymbol{z})$ is injected into $\phi_{i}$. Subsequently, the original word $x_{i}$ is replaced with a word $\tilde{x}_{i}$ whose embedding $\phi(\tilde{x}_{i})$ is closest to the $\tilde{\phi}_{i}$ within the embedding space. The previous framework of the word-level DP satisfies $\epsilon{d}$-MetricDP, where $d:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}_{+}$ is a distance defined within sentence space (e.g., in Euclidean metric space, $d(\boldsymbol{x},\boldsymbol{x}^{\prime})=\sum_{i=1}^{n}\|\phi(x_{i})-\phi(x^{\prime}_{i})\|$ for $\boldsymbol{x}$ and $\boldsymbol{x}^{\prime}$ with the same length). The outline of the word-level DP is provided in Algorithm 2.
Algorithm 2 Word-level DP [24].
3:for ${i}\leftarrow{1},{2},\dots,{n}$ do
4:     Obtain embedding $\phi_{i}\leftarrow\phi(x_{i})$
5:     Sample noise $\boldsymbol{z}\sim{p}_{\epsilon}(\boldsymbol{z})$
6:     Perturb embedding $\tilde{\phi}_{i}\leftarrow\phi_{i}+\boldsymbol{z}$
Sentence-level DP approaches. Compared to the word-level DP, the sentence-level DP aims to achieve DP at the full document rather than a single word. A common approach is to leverage LMs for paraphrasing tasks to achieve DP [41, 57]. Given a private text $\boldsymbol{x}$, the prompt for paraphrasing is constructed by prompt template $\boldsymbol{p}_{T}$ and $\boldsymbol{x}$. In an autoregressive language model, the output logit in the last layer of the decoder is denoted as $\boldsymbol{u}\in\mathbb{R}^{|\mathcal{V}|}$ when modeling the next token. To limit the sensitivity, the logit is clipped as $\bar{\boldsymbol{u}}=\textsc{Clip}(\boldsymbol{u},C)$, where the clipping function is $\textsc{Clip}(\boldsymbol{u},C)=\boldsymbol{u}\cdot\min(1,C/\|\boldsymbol{u}\|)$. The probability for each word in the vocabulary is computed as $P_{j}=\textsc{Softmax}(\bar{\boldsymbol{u}})={\exp(\bar{\boldsymbol{u}}_{j}/T)}/{\sum_{j=1}^{|\mathcal{V}|}\exp(\bar{\boldsymbol{u}}_{j}/T)}$, where $T$ represents the temperature for controlling randomness level of output. The next token is sampled from the conditional probability distribution $P$ over the vocabulary $\mathcal{V}$. Note that the sampling procedure can be viewed as the exponential mechanism [42], and the temperature is related to the privacy budget. Given the output length $m$, the clipping constant $C$, the paraphrasing procedure is $2{m}{C}{\epsilon}/{T}$-LDP following the exponential mechanism. The outline of the sentence-level method is presented in Algorithm 3.
Algorithm 3 Sentence-level DP (based on paraphrasing) [41, 32].
2:sanitized text $\tilde{\boldsymbol{x}}$
4:$\boldsymbol{x}^{\prime}\leftarrow\emptyset$
5:for ${i}\leftarrow{1},{2}\dots,{m}$ do
6:     Compute logit $\boldsymbol{u}$ with LM using prompt $\boldsymbol{y}$
9:     Sample the next token ${v}$ from $\mathcal{V}$ using $\mathbb{P}$
Comparison between the word-level DP and the sentence-level DP.
For texts with different lengths, the word-level DP can not provide privacy guarantees for two texts of the same length because it outputs the fixed-length text. Nevertheless, the sentence-level DP can adapt to texts of varying lengths, with a linear increase in privacy budget as the length of the output text grows. Table 1 lists the summary of differentially private text sanitization methods. For the privacy definition, based on metric DP, the distance used in the word-level DP mainly reflects word differences within the text but ignores semantic differences. In contrast, the sentence-level DP follows the Local DP (LDP), where any two texts are adjacent, making it hard to achieve a practical privacy budget. For the implementation of DP, the word-level DP introduces perturbation at the word level, and the sentence-level DP perturbs the semantic information of the text. Note that with a small privacy budget, the word-level DP adds a large-scale noise to the embeddings of the original words, significantly changing the word types and resulting in low utility.
Appendix D PII Classes
Table 15 gives the semantics and examples of PII classes.
Table 15: Semantics and Examples of PII Classes.
Name Semantics Examples
cardinal cardinal value 181
date date value 2014
event event name Hurricane Ana
fac building name the Stade de France
gpe geo-political entity Egypt
language language name Spanish
law law name The Local Audit and Accountability Act 2014
loc location name the Central Pacific
money money name €375
norp affiliation Conservatives
ordinal ordinal value second
org organization name Muslim Brotherhood
percent percent value 42.1%
person person name Mohamed Badie
product product name Ziyuan I-04
quantity quantity name 82 metres
time time value 2:02:57
work of art name of work of art The BRW Rich 200, 2014
Figure 3: Figure 3, Figure 3 and Figure 3 provide the correlation between the Succ, Recall, Precision and privacy budgets, respectively. Figure 3 shows the Precision and Recall for different PII classes on WikiMIA using word-level DP under $\epsilon=8$.
@@ Appendix D PII Classes
Table 16: Results of the white-box fine-tuning-based attacks against the word-level DP on WikiMIA.
@@
Appendix E Large Language Models Used in the Experiments

### v3

@@ V-B Datasets
TABLE X: Results of the black-box instruction-based attacks against the word-level DP on WikiMIA.
@@
TABLE XI: Results of the black-box instruction-based attacks against the sentence-level DP on WikiMIA.
@@
TABLE XII: Results of the white-box fine-tuning-based attacks against the word-level DP on Pile-$C$.
@@
TABLE XIII: Results of the white-box fine-tuning-based attacks against the word-level DP on Pile-Wiki.
@@
TABLE XIV: Results of the white-box fine-tuning-based attacks against the word-level DP on Pile-Enron.
@@
TABLE XV: Example of the white-box fine-tuning attacks against the word-level DP on Pile-CC and Pile-Enron.
Pile-$C$ Pile-Enron
@@
V-C Differentially Private Text Sanitization Approaches
In the experiments, we choose MadLib [25] (Alg. 1) as the word-level DP in our experiments, since this is the commonly used DP word-level implementation [41, 20]. In MadLib, the distance function is defined in Euclidean space and Laplace noise is added to the word embedding vector. We choose DP-Prompt [57] (Alg. 2) as the sentence-level DP in our experiments, since Utpala et al. [57] demonstrate the DP-Prompt achieves better utility compared to other approaches when using ChatGPT. In DP-Prompt, we consider using ChatGPT-3.5 as base model to generate the paraphrased text.
V-D Metrics
@@
  • Recall and Precision. In our study, we focus on the privacy leakage of sensitive content in the text rather than the shared vocabulary (e.g., time, addresses, and names), which are more prone to causing privacy concerns; whereas common words like “the” and “this” do not lead to privacy issues. That is, we aim to recover the privacy-sensitive content from the original text. Hence, we extract and mark PII sequences (see Sec. II for details) from the text and consider these sequences as the privacy within the given text. Formally, we denote the set of PII sequences in the original text $\boldsymbol{x}$ as $C$, in the DP-sanitized text $\tilde{\boldsymbol{x}}$ as $\tilde{C}$, and in the reconstructed text $\hat{\boldsymbol{x}}$ as $\hat{C}$. We define recall and precision for the reconstruction attacks below.
@@ V-E Implementation
To tag PII sequences from text, we leverage Flair [10] as the PII extraction model released in HuggingFace [2]. Specifically, Flair is a NER framework based on Transformer networks to classify tokens as PII sequences. It is noteworthy that in our experiments we did not store any PII as auxiliary knowledge for our attacks. For the models giving weight access (e.g. LLaMA, Gemma and OPT), we downloaded checkpoints from HuggingFace. For the models not giving weight access (e.g. ChatGPT, Claude), we query models via black-box APIs available online (e.g., OpenAI API [4] and Anthropic AI API [1]). For DP text sanitization approaches, we implement both the word-level DP and the sentence-level DP based on DP-Prompt repository [56]. Note that we implement the word-level DP using 50-dimensional GloVe [49] vectors (following existing works [41, 57]) as word embeddings. We implement our reconstruction attacks based on PyTorch [47] and Transformers [61]. To fine-tune the pre-trained LLMs, we leverage LoRA [29] based on PEFT [39] library, a parameter-efficient fine-tuning method, to reduce the trainable parameters due to the limited computational resource. To obtain the sensitive information similarity by LLMs, we employ ChatGPT-3.5 as the base model. Tab. VII presents the prompt templates used for evaluating the sensitive information similarity. Finally, we run all the experiments on a GPU cluster, which is equipped with NVIDIA RTX3090, RTX4090 and A800.
V-F Results of Black-box Instruction-based Attacks
@@
Setup. To evaluate the effectiveness of the black-box instruction-based attack, we conduct experiments on WikiMIA using both the word-level DP and the sentence-level DP under various privacy budgets ($\epsilon=\{4,8,12\}$ for the word-level DP [24] or $T=\{2.0,1.5,1.0\}$ for the sentence level DP [57], where the relationship between $T$ and the privacy budget can be found in Sec. II) with instruction-tuned LLMs, including LLaMA, Gemma, ChatGPT, and Claude. We adjust the prompt per Fig. 2 for different models and query models with the sanitized text.
@@
Results. Tab. VIII and Tab. IX provides an example of the black-box instruction-based attack on WikiMIA under $\epsilon=8$ using ChatGPT-4, Claude-3-Opus, LLaMA-3-70B and Gemma-2-27B for the word-level DP and the sentence-level DP, repectively. The original text $\boldsymbol{x}$, sanitized text $\tilde{\boldsymbol{x}}$ and reconstructed text $\hat{\boldsymbol{x}}$ are listed, where the highlighted text is PII sequences extracted by the NER model, and the text in the red box contains the PII sequences that remain unchanged. For word-level DP, we observe that the reconstructed text can correct the errors in the sanitized text (e.g., reconstructed text using ChatGPT-4 correct “31st” to “51st”). Compared to word-level DP, sentence-level DP retains most of the PII sequences from the original text, thereby fewer PII sequences are successfully reconstructed.
@@
The numeric results on WikiMIA for the word-level DP are detailed in Tab. X and Tab. XI, respectively. For the word-level DP, in a practical privacy budgets (e.g., $\epsilon=8$ and $\epsilon=12$), the black-box instruction-based attacks successfully reconstruct sensitive information at the word level with high probability. Nevertheless, the black-box instruction-based attacks have no significant effect under a smaller privacy budget due to substantial discrepancies between the original and sanitized texts. For the sentence-level DP, we observe that Succ, Recall, and Precision are lower than the word-level DP, but Score is higher. The reason behind it is in two folds. First, the sentence-level DP fails to remove the sensitive information (see examples as the text enclosed in the red boxes in Tab. XI). Second, the metrics of recall, precision, and succ measures the difference between the DP-sanitized text and the reconstructed text. Hence, these three metrics do not show a promising results.
@@
V-G Results of White-box Fine-tuning-based Attacks
@@
Results. Tab. XV provides an example of the white-box fine-tuning-based attacks on Pile-CC under $\epsilon=8$ with Pythia-6.9B, OPT-6.7B and GPT-J-6B for the word-level DP. We observe that, the open-source model reconstructs few PII sequences compared to close-source models (e.g., ChatGPT, Claude).
@@
Tab. XII, Tab. XIII, and Tab. XIV report the reconstruction results and Precision on Pile-CC, Pile-Wiki and Pile-Enron, respectively, where the metrics include Succ, Recall and Precision. Similar to the black-box instruction-based attacks, we observe that the model with more parameters can achieve higher performance metrics. For Pile-CC, the metrics including Succ, Recall, and Precision are lower than Pile-Wiki and Pile-Entron, this may be because Pile-CC is not composed of a specific type of data. Therefore, the model cannot learn from a specific data distribution like fact (Pile-Wiki) or email (Pile-Enron) distribution.
V-H Ablation Studies
The original text of the DP-sanitized one was not seen by LLMs. Tab. XVI shows the results of prompting fine-tuned white-box LLMs with DP-sanitized text, where the original text was not seen by the LLMs. From the results, we can see that the LLMs we used still achieve acceptable recovery outcome compared to the cases where the original text of the DP-sanitized ones was seen by LLMs.
@@ V-H Ablation Studies
TABLE XVI: Results of the white-box fine-tuning-based attacks against the word-level DP on WikiMIA.
@@
Privacy budgets. In this experiment, we aim to investigate how privacy budget affects the quality of reconstruction attacks. Intuitively, privacy budgets control the level of modification for the DP text sanitization approach. For experimental settings, we vary the privacy budget from 1 to 32 and apply the black-box instruction-based attack to LLaMA models. Fig. 3, Fig. 3, Fig. 3 respectively provide the correlation between the Succ, Recall, Precision, and the privacy budget. We can observe that, for a greater privacy budget ($\epsilon\geq 8$), our attacks have higher precision and recall, indicating that our approach can attack the DP text sanitization approach under a practical privacy budget. In contrast, for a smaller privacy budget, both precision and recall are low, indicating that only a minimal amount of sensitive information can be recovered.
Figure 3: Fig. 3, Fig. 3 and Fig. 3 provide the correlation between the Succ, Recall, Precision and privacy budgets, respectively. Fig. 3 shows the Precision and Recall for different PII classes on WikiMIA using word-level DP under $\epsilon=8$.
VI Related Work
Comparision with existing studies. Existing works discovered the ”current” privacy concerns, which could be addresses by DP, the de facto privacy standard. However, our work, focusing on ”future” privacy concerns as motivated in Sec. I, demonstrates that DP might have the potential vulnerability in the era of LLMs, which can recover the privacy from the DP-sanitized text.
VII Mitigation & Discussion
This section gives potential mitigation strategies to further preserve privacy against the privacy leakage discovered in this paper. We also discussed the technical limitation of this paper. The ethics statement is attached in Appx. VIII-D.
PII classes. We also conducted experiments to show the performance of our attacks for different PII classes, which have different degrees of sensitive information leakage. Direct identifiers (e.g., names) leak more sensitive information than quasi-identifiers (e.g., addresses). For experimental settings, we select six classes (e.g., person, address, organization), and apply the black-box instruction-based attack to ChatGPT-4 with the word-level DP under privacy budget $\epsilon=8.0$. Fig. 3 in Sec. V-H shows the precision and recall for different PII classes on WikiMIA. The results demonstrate that our reconstruction attacks exhibit varying levels of effectiveness in recovering different types of sensitive information.
Limitation. In this study, we could not conduct useful ablation studies to determine whether the training data played a crucial role in LLMs’ ability to recover privacy. On one hand, performing such ablation studies on black-box attacks is infeasible due to the unpublished training data of the LLMs, despite their promising recovery rates. On the other hand, for LLMs with specified training data, we tested white-box attacks using DP-sanitized text, where the corresponding original text was not included in the LLMs’ training data. Specifically, we prompted the same fine-tuned LLMs as those in Tab. XIII using DP-sanitized text from WikiMIA (which was guaranteed not to have been seen by the LLMs). The results, shown in Tab. XVI in Sec. V-H, indicate similar recovery rates to those in Tab. XIII. However, these similar results further confirm that white-box LLMs trained on a limited number of text data do not exhibit sufficient capability to perform our attacks.
VIII Conclusion
Appendix
VIII-A Language Models
VIII-B Word- and Sentence-level DP Approaches
Word-level DP approaches. The goal of the word-level DP is to perturb each word in a sentence under the metric DP. Formally, consider a sequence $\boldsymbol{x}=\{x_{1},x_{2},\dots,x_{n}\}$ with $n$ tokens, each token $x_{i}$ is convert to a $m$-dimensional word embedding $\phi_{i}=\phi(x_{i})$ by a pre-trained word embedding model $\phi:\mathcal{V}\rightarrow\mathbb{R}^{m}$, where $\mathcal{V}$ represents the vocabulary. To obtain the noisy embedding $\tilde{\phi}_{i}$, a noise $\boldsymbol{z}$ sampled from a multivariate probability distribution $p_{\epsilon}(\boldsymbol{z})$ is injected into $\phi_{i}$. Subsequently, the original word $x_{i}$ is replaced with a word $\tilde{x}_{i}$ whose embedding $\phi(\tilde{x}_{i})$ is closest to the $\tilde{\phi}_{i}$ within the embedding space. The previous framework of the word-level DP satisfies $\epsilon{d}$-MetricDP, where $d:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}_{+}$ is a distance defined within sentence space (e.g., in Euclidean metric space, $d(\boldsymbol{x},\boldsymbol{x}^{\prime})=\sum_{i=1}^{n}\|\phi(x_{i})-\phi(x^{\prime}_{i})\|$ for $\boldsymbol{x}$ and $\boldsymbol{x}^{\prime}$ with the same length).
Sentence-level DP approaches. Compared to the word-level DP, the sentence-level DP aims to achieve DP at the full document rather than a single word. A common approach is to leverage LMs for paraphrasing tasks to achieve DP [41, 57]. Given a private text $\boldsymbol{x}$, the prompt for paraphrasing is constructed by prompt template $\boldsymbol{p}_{T}$ and $\boldsymbol{x}$. In an autoregressive language model, the output logit in the last layer of the decoder is denoted as $\boldsymbol{u}\in\mathbb{R}^{|\mathcal{V}|}$ when modeling the next token. To limit the sensitivity, the logit is clipped as $\bar{\boldsymbol{u}}=\textsc{Clip}(\boldsymbol{u},C)$, where the clipping function is $\textsc{Clip}(\boldsymbol{u},C)=\boldsymbol{u}\cdot\min(1,C/\|\boldsymbol{u}\|)$. The probability for each word in the vocabulary is computed as $P_{j}=\textsc{Softmax}(\bar{\boldsymbol{u}})={\exp(\bar{\boldsymbol{u}}_{j}/T)}/{\sum_{j=1}^{|\mathcal{V}|}\exp(\bar{\boldsymbol{u}}_{j}/T)}$, where $T$ represents the temperature for controlling randomness level of output. The next token is sampled from the conditional probability distribution $P$ over the vocabulary $\mathcal{V}$. Note that the sampling procedure can be viewed as the exponential mechanism [42], and the temperature is related to the privacy budget. Given the output length $m$, the clipping constant $C$, the paraphrasing procedure is $2{m}{C}{\epsilon}/{T}$-LDP following the exponential mechanism.
VIII-C Large Language Models Used in the Experiments